### v1

@@ 3.1 Forecasting-Specific Meta-Knowledge
Search Space Constraints.
@@
Few-Shot Task Context (Optional).
When available, a small number of representative input–output examples may be included to provide additional contextual grounding. These examples help the LLM reason about scale, variability, and noise characteristics of the data without relying on full dataset access.
Target Loss Criterion.
The optimization process is governed by a target loss threshold derived from the BO phase. Specifically, the target loss is defined as
$\text{Target\_Loss}=L_{\min}^{\text{BO}}-\varepsilon,$
where $L_{\min}^{\text{BO}}$ is the best loss observed during BO and $\varepsilon>0$ is a user-defined margin. This criterion serves as a stopping condition for the iterative LLM-guided optimization, ensuring that computational effort is focused on refining strong configurations while avoiding unnecessary evaluations.
@@ 3.2 Structured Prompt Design
The meta-knowledge is embedded within a structured framework, zero-shot prompt that conditions the LLM to generate hyperparameter updates. The prompt is designed to balance expressive reasoning with strict structural constraints. It enforces a single valid JSON output, prohibits duplicate hyperparameters, and requires adherence to predefined types and search space bounds. Internal reasoning fields encourage the LLM to justify each recommendation based on the provided meta-knowledge, improving transparency without exposing free-form text that could destabilize downstream processing. This design minimizes hallucinations, ensures consistency across iterations, and enables reliable integration of LLM outputs into the optimization loop.
@@ 3.3 LLM-AutoOpt Optimization Workflow
Figure 1 illustrates the overall optimization workflow, which consists of two tightly coupled phases.
Figure 1: Automated hyperparameter optimization guided by LLM reasoning.
Phase 1: Meta-Knowledge Construction.
@@
Phase 2: LLM-Guided Iterative Refinement.
@@
By integrating Bayesian initialization with LLM-based in-context reasoning, LLM-AutoOpt enables sample-efficient, interpretable, and forecasting-aware hyperparameter optimization. Unlike conventional HPO methods that operate on isolated trials, the framework allows the LLM to reason over historical optimization behavior and task-specific structure, leading to more informed and adaptive optimization trajectories.
@@ 4.1.1 Dataset and Preprocessing
We conduct experiments on the Jena Climate multivariate time-series dataset, collected at the meteorological station of the Max Planck Institute for Biogeochemistry in Jena, Germany. The dataset contains high-frequency measurements recorded every 10 minutes between January 1, 2004 and December 31, 2020 and is widely used as a benchmark for multivariate forecasting tasks [24]. Each timestamp includes 14 meteorological variables such as air temperature, humidity, atmospheric pressure, and wind characteristics. To ensure computational tractability while preserving temporal structure, we extract a continuous six-month subset of the data. The CO2 variable is excluded due to extensive missing values. The remaining variables—air temperature (T), relative humidity (rh), atmospheric pressure (p), and wind speed (wv)—are selected for their relevance and complementary physical interpretations.
@@ 4.1.2 Forecasting Model
Bidirectional LSTM (Bi-LSTM).
We select the Bidirectional Long Short-Term Memory (Bi-LSTM) network as the candidate forecasting model due to its widespread use and effectiveness in multivariate time-series modeling [25]. Bi-LSTMs process sequences in both forward and backward directions, enabling the capture of long-range temporal dependencies.
Optimized Hyperparameters.
@@
Batch size Number of samples per update; affects training stability and generalization
Hidden size Number of LSTM units; larger size captures complex patterns but may overfit
Number of layers Stacked LSTM layers; more layers capture long-term dependencies but may overfit
Dropout Fraction of neurons ignored during training; prevents overfitting, improves robustness
Learning rate (lr) Step size for weight updates; high lr speeds up training but may be unstable
Epochs Number of passes over data; more epochs improve learning but may overfit
Lag Number of previous time steps as input; captures temporal dependencies
Optimizer Method for updating weights (e.g., Adam, SGD); affects convergence and accuracy
@@ 4.1.3 Large Language Model Backbone
The LLM backbone used for hyperparameter reasoning is Qwen2.5-72B, a large-scale language model developed by the Qwen team at Alibaba Group. The model is pretrained on multilingual and multimodal data and further aligned using high-quality instruction datasets. Its strong contextual reasoning capabilities make it well-suited for structured decision-making tasks such as hyperparameter optimization. Prior studies show that Qwen2.5-72B achieves low regret as the number of support tasks increases, outperforming alternative LLMs in optimization-oriented settings [22, 23]. All LLM outputs are validated against predefined constraints to ensure correctness and feasibility. To balance deterministic behavior and mitigate hallucinations, the decoding temperature parameter was set to 0.2.
@@ 4.1.4 Meta-Knowledge Prompt Configuration
Data Meta-Knowledge.
@@
Table 3: Meta-features used to describe the input data
Meta-feature Description
count Number of observations in the column
missing Number of missing (NaN) values
mean Average value of the data
std Standard deviation of the data
min Minimum value
25% 25th percentile
50%_median Median (50th percentile)
75% 75th percentile
max Maximum value
skewness Measure of asymmetry of the distribution
kurtosis Measure of "tailedness" of the distribution
range Difference between max and min
iqr Interquartile range (75%-25%)
variance Variance of the data
acf_lag_1,3,6,12,24 Autocorrelation at specified lags
pacf_lag_1,3,6,12,24 Partial autocorrelation at specified lags
trend_strength Absolute slope of linear trend (regression)
adf_stat ADF test statistic for stationarity
adf_pvalue p-value of ADF test
num_peaks Number of local maxima
num_troughs Number of local minima
zero_ratio Fraction of zero values in series
outlier_ratio Fraction of outliers based on 1.5*IQR rule
coef_of_variation Relative variability (std/mean)
nonlinearity_proxy Std of first differences, indicates nonlinearity
trend_strength_decomp Strength of trend component from decomposition
seasonal_strength_decomp Strength of seasonal component from decomposition
residual_strength Variance of residual component from decomposition
Model Meta-Knowledge.
@@
Table 4: Historical Bayesian Optimization trials used as contextual guidance for LLM-based hyperparameter recommendation
Trial ID Lag Hidden Size Layers Dropout LR Batch Epochs Optimizer
@@
Target Loss Criterion.
The optimization loop terminates when the RMSE reaches a target threshold defined as the best BO loss minus a margin $\varepsilon=0.09$, corresponding to a target RMSE of 1.00.
Prompt Constraints.
Strict output constraints enforce valid JSON formatting, prohibit repeated hyperparameters, and ensure compliance with search space boundaries. These constraints minimize hallucination and ensure that recommendations are directly executable.
4.1.5 Baselines and Method Variant
We evaluate LLM-AutoOpt against several baselines. First, we consider a baseline Bi-LSTM model with randomly initialized hyperparameters, denoted as Bi-LSTM-Base, using the configuration specified in Table 2. Second, we compare against a Bi-LSTM whose hyperparameters are selected via Bayesian Optimization, referred to as Bi-LSTM-BO. Finally, we assess an ablated variant of our method that excludes meta-knowledge from the prompt, denoted as LLM-AutoOpt-NoMeta, to isolate the impact of meta-knowledge on optimization performance.
@@ 4.1.6 Evaluation Metrics
We evaluate performance using three complementary metrics:
Forecasting Accuracy (RMSE).
Root Mean Square Error (RMSE) is used to assess predictive accuracy. It penalizes large errors and is well suited for continuous forecasting tasks. RMSE serves both as the evaluation metric and as feedback for the LLM-guided optimization loop.
Training Time.
Training time measures the computational cost of fitting the model given a fixed hyperparameter configuration. This metric captures the downstream cost induced by different optimization strategies.
Optimization Time.
Optimization time measures the total wall-clock time required to complete the hyperparameter optimization process, including both BO and LLM-guided iterations.
4.2 Results
@@ 4.2.1 RQ1: Predictive Performance
Figure 2: Validation RMSE convergence across optimization strategies
Table 5: Comparison of Mean RMSE, Average Training Time, and Optimization Time for Different Methods (with Standard Deviations)
@@ 4.2.4 RQ4: Impact of Meta-Knowledge on Optimization Quality
Table 6: Hyperparameter configurations and RMSE results across optimization trials with method comparisons
@@
4.2.5 RQ5: Interpretability of LLM-Guided Optimization
@@
4.3 Discussion
The experimental results demonstrate that incorporating meta-knowledge into LLM-based hyperparameter optimization yields consistent improvements in forecasting accuracy and optimization behavior. Across all evaluated settings, LLM-AutoOpt achieves lower RMSE and faster convergence compared to manually initialized models, Bayesian Optimization, and LLM-based optimization without meta-knowledge. These gains are reflected not only in quantitative metrics but also in the qualitative forecasting results, where meta-informed configurations produce smoother and more accurate predictions that better capture temporal dynamics.
From a methodological perspective, the results indicate that meta-knowledge plays a dual role. First, it improves optimization effectiveness by providing the LLM with structured priors over the data distribution, model architecture, and historical optimization outcomes. This reduces unproductive exploration and allows the LLM to focus on meaningful hyperparameter adjustments, such as tuning lag length, regularization strength, or network depth. Second, meta-knowledge stabilizes the reasoning process itself. In contrast to the LLM-AutoOpt-NoMeta setting, where the LLM often relies on generic heuristics (e.g., increasing epochs or model size), LLM-AutoOpt generates decisions that are explicitly grounded in dataset characteristics and model constraints.
Interpretability emerges as a key advantage of the proposed framework. The reasoning traces produced by LLM-AutoOpt reveal how specific meta-features and architectural considerations influence hyperparameter choices, enabling a transparent link between data properties and optimization decisions. This stands in contrast to Bayesian Optimization, which, while computationally efficient, offers limited insight into why particular configurations are preferred. As such, LLM-AutoOpt provides a more explainable alternative for practitioners seeking both performance and interpretability.
However, these benefits come at the cost of computational efficiency. BO remains substantially faster in terms of both optimization and training time, highlighting a trade-off between speed and interpretability. This suggests that LLM-AutoOpt is best suited for scenarios where accuracy, robustness, and explainability are prioritized over raw optimization speed. Additionally, the framework’s performance depends on the quality of prompt design and the relevance of extracted meta-features, indicating that careful engineering is required to fully realize its benefits.
Finally, while the results are encouraging, they are based on a single multivariate forecasting benchmark. Further evaluation across diverse datasets, architectures, and forecasting horizons is necessary to assess the generality and scalability of the approach.

### v2

@@ 3.3.3 Model Specification
3.3.4 Search Space Constraints
@@
3.3.5 Few-Shot Context (Optional)
Representative input-output examples may be included to provide additional grounding, allowing the LLM to reason about scale, variability, and noise characteristics.
3.3.6 Target Loss Criterion
The iterative optimization stops once the target loss threshold is achieved:
$\mathcal{L}_{\text{target}}=\mathcal{L}_{\text{BO}}^{\min}-\epsilon,\quad\epsilon>0$ (3)
@@ 3.4 Structured Prompt Design
The meta-knowledge is embedded within a structured framework, a zero-shot prompt that conditions the LLM to generate hyperparameter updates. The prompt is designed to balance expressive reasoning with strict structural constraints. It enforces a single valid JSON output, prohibits duplicate hyperparameters, and requires adherence to predefined types and search space bounds. Internal reasoning fields encourage the LLM to justify each recommendation based on the provided meta-knowledge, improving transparency without exposing free-form text that could destabilize downstream processing. This design minimizes hallucinations, ensures consistency across iterations, and enables reliable integration of LLM outputs into the optimization loop. This ensures interpretability, minimizes hallucinations, and guarantees consistency for downstream training.
@@ 3.5 LLM-AutoOpt Optimization Workflow
The workflow consists of two main phases as shown in Figure 1:
Phase 1: Meta-Knowledge Construction
@@
Phase 2: LLM-Guided Iterative Refinement
@@
Advantages: The framework is sample-efficient, interpretable, task-aware, and provides iterative refinement that adapts to historical optimization behavior and dataset characteristics.
Figure 1: Automated hyperparameter optimization workflow guided by LLM reasoning.
@@ 4.1.1 Dataset and Preprocessing
We conduct experiments on the Jena Climate multivariate time-series dataset, collected at the meteorological station of the Max Planck Institute for Biogeochemistry in Jena, Germany. The dataset contains high-frequency measurements recorded every 10 minutes between January 1, 2004 and December 31, 2020 and is widely used as a benchmark for multivariate forecasting tasks [bib34]. Each timestamp includes 14 meteorological variables such as air temperature, humidity, atmospheric pressure, and wind characteristics.To ensure computational tractability while preserving temporal structure, we extract a continuous six-month subset of the data. The CO2 variable is excluded due to extensive missing values. The remaining variables, air temperature (T), relative humidity (rh), atmospheric pressure (p), and wind speed (wv), are selected due to their relevance and complementary physical interpretations.
@@ 4.1.2 Forecasting Model
Bidirectional LSTM (Bi-LSTM)
We select the Bidirectional Long Short-Term Memory (Bi-LSTM) network as the candidate forecasting model due to its widespread use and effectiveness in multivariate time-series modeling [bib35]. Bi-LSTMs process sequences in both forward and backward directions, enabling the capture of long-range temporal dependencies.
Optimized Hyperparameters
@@
Batch size Number of samples per update; affects training stability and generalization.
Hidden size Number of LSTM units; larger size captures complex patterns but may overfit.
Number of layers Stacked LSTM layers; more layers capture long-term dependencies but may overfit.
Dropout Fraction of neurons ignored during training; prevents overfitting, improves robustness.
Learning rate (lr) Step size for weight updates; high lr speeds up training but may be unstable.
Epochs Number of passes over data; more epochs improve learning but may overfit.
Lag Number of previous time steps used as input; captures temporal dependencies.
Optimizer Method for updating weights (e.g., Adam, SGD); affects convergence and accuracy.
@@ 4.1.3 Large Language Model Backbone
The LLM backbone used for hyperparameter reasoning is Qwen2.5-72B, a large-scale language model developed by the Qwen team at Alibaba Group. The model is pretrained on multilingual and multimodal data and further aligned using high-quality instruction datasets. Its strong contextual reasoning capabilities make it well-suited for structured decision-making tasks such as hyperparameter optimization. Prior studies show that Qwen2.5-72B achieves low regret as the number of support tasks increases, outperforming alternative LLMs in optimization-oriented settings [bib17, bib18]. All LLM outputs are validated against predefined constraints to ensure correctness and feasibility. To balance deterministic behavior and mitigate hallucinations, the decoding temperature parameter was set to 0.2.
@@ 4.1.4 Meta-Knowledge Prompt Configuration
Data Meta-Knowledge
@@
Table 3: Meta-feature definitions and formulas used to summarize statistical and temporal properties of a univariate time series $\{x_{t}\}_{t=1}^{T}$.
Meta-feature Definition / Formula
count $n=\sum_{t=1}^{T}1$ (number of observed elements)
missing $\mathrm{missing}=\sum_{t=1}^{T}I(\text{NA}(x_{t}))$
mean $\bar{x}=\dfrac{1}{n}\sum_{t=1}^{n}x_{t}$
std $s=\sqrt{\dfrac{1}{n-1}\sum_{t=1}^{n}(x_{t}-\bar{x})^{2}}$
min $\min_{t}x_{t}$
25% first quartile $Q_{1}=x_{(0.25)}$ (25th percentile)
50%_median median $\mathrm{median}=x_{(0.5)}$ (50th percentile)
75% third quartile $Q_{3}=x_{(0.75)}$ (75th percentile)
max $\max_{t}x_{t}$
skewness $\text{skewness}=\frac{\frac{1}{n}\sum_{t=1}^{n}(x_{t}-\bar{x})^{3}}{\left(\frac{1}{n}\sum_{t=1}^{n}(x_{t}-\bar{x})^{2}\right)^{3/2}}$
kurtosis $\text{kurtosis}=\frac{\frac{1}{n}\sum_{t=1}^{n}(x_{t}-\bar{x})^{4}}{\left(\frac{1}{n}\sum_{t=1}^{n}(x_{t}-\bar{x})^{2}\right)^{2}}-3$ (excess kurtosis)
range $\mathrm{range}=\max_{t}x_{t}-\min_{t}x_{t}$
iqr $\mathrm{IQR}=Q_{3}-Q_{1}$
variance $\mathrm{Var}(x)=s^{2}$
acf_lag_k $\rho_{k}=\frac{\sum_{t=k+1}^{T}(x_{t}-\bar{x})(x_{t-k}-\bar{x})}{\sum_{t=1}^{T}(x_{t}-\bar{x})^{2}},\qquad k\in\{1,3,6,12,24\}$
pacf_lag_k Partial autocorrelation at lag $k$, $\phi_{kk}$ from the Yule–Walker or Levinson–Durbin equations (PACF values at $k\in\{1,3,6,12,24\}$).
trend_strength Fit linear regression $x_{t}=a+bt+e_{t}$, then $\text{trend\_strength}=|b|$ (absolute slope)
adf_stat ADF test regression: $\Delta x_{t}=\alpha+\beta t+\gamma x_{t-1}+\sum_{i=1}^{p}\delta_{i}\Delta x_{t-i}+e_{t}$. ADF statistic = $t$-statistic of $\hat{\gamma}$.
adf_pvalue p-value associated with the ADF statistic (computed via the Dickey–Fuller distribution).
num_peaks $\sum_{t=2}^{T-1}I(x_{t-1}<x_{t}\wedge x_{t}>x_{t+1})$ (local maxima count)
num_troughs $\sum_{t=2}^{T-1}I(x_{t-1}>x_{t}\wedge x_{t}<x_{t+1})$ (local minima count)
zero_ratio $\dfrac{1}{n}\sum_{t=1}^{n}I(x_{t}=0)$
outlier_ratio Fraction outside Tukey fences: $\dfrac{1}{n}\sum_{t=1}^{n}I\big(x_{t}<Q_{1}-1.5\cdot\mathrm{IQR}\ \text{or}\ x_{t}>Q_{3}+1.5\cdot\mathrm{IQR}\big)$
coef_of_variation $\mathrm{CV}=\dfrac{s}{|\bar{x}|}$
nonlinearity_proxy $\mathrm{std}(\Delta x_{t})=\sqrt{\mathrm{Var}(x_{t}-x_{t-1})}$ (std of first differences)
trend_strength_decomp From decomposition $x_{t}=T_{t}+S_{t}+R_{t}$: $\mathrm{trend\_strength\_decomp}=\dfrac{\mathrm{Var}(T_{t})}{\mathrm{Var}(x_{t})}$
seasonal_strength_decomp $\mathrm{seasonal\_strength\_decomp}=\dfrac{\mathrm{Var}(S_{t})}{\mathrm{Var}(x_{t})}$
residual_strength $\mathrm{residual\_strength}=\dfrac{\mathrm{Var}(R_{t})}{\mathrm{Var}(x_{t})}$
Model Meta-Knowledge
@@
Table 4: Historical Bayesian Optimization trials used as contextual guidance for LLM-based hyperparameter recommendation.
Trial ID Lag Hidden Layers Dropout LR Batch Epochs Optimizer
@@ 4.1.5 Evaluation Metrics
We evaluate performance using three complementary metrics designed to assess both solution quality and computational efficiency:
Forecasting Accuracy (RMSE)
Root Mean Square Error (RMSE) is used to assess predictive accuracy. It is defined as:
$\text{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}$ (4)
where $y_{i}$ is the true value at step $i$ and $\hat{y}_{i}$ is the predicted value, with $N$ being the total number of test samples. RMSE penalizes large errors quadratically and is well suited for continuous forecasting tasks. RMSE serves both as the evaluation metric and as feedback signal for the LLM-guided optimization loop, allowing the framework to iteratively refine hyperparameter configurations based on validation performance.
Training Time
Training time $T_{\text{train}}(\lambda)$ measures the computational cost of fitting the model given a fixed hyperparameter configuration:
$T_{\text{train}}(\lambda)=t_{\text{end}}-t_{\text{start}}$ (5)
where $t_{\text{start}}$ and $t_{\text{end}}$ are the wall-clock timestamps marking the beginning and end of model training. This metric captures the downstream computational cost induced by different optimization strategies. Hyperparameters such as batch size, number of epochs, and hidden dimensions directly influence training time, making this metric crucial for evaluating practical efficiency.
Optimization Time
Optimization time measures the total wall-clock time required to complete the hyperparameter optimization process. We compute optimization time separately for each technique:
Bayesian Optimization Time:
$T_{\text{opt}}^{\text{BO}}=\sum_{i=1}^{N_{\text{BO}}}\left(T_{\text{acq}}^{(i)}+T_{\text{train}}^{(i)}+T_{\text{eval}}^{(i)}\right)$ (6)
where $N_{\text{BO}}$ is the number of BO iterations, $T_{\text{acq}}^{(i)}$ is the time to compute the acquisition function and select the next candidate, $T_{\text{train}}^{(i)}$ is the training time for configuration $i$, and $T_{\text{eval}}^{(i)}$ is the validation evaluation time.
LLM-Guided Optimization Time:
$T_{\text{opt}}^{\text{LLM}}=\sum_{t=1}^{M}\left(T_{\text{LLM}}^{(t)}+T_{\text{train}}^{(t)}+T_{\text{eval}}^{(t)}\right)$ (7)
where $M$ is the number of LLM-guided refinement iterations, $T_{\text{LLM}}^{(t)}$ is the time for the LLM to generate a hyperparameter recommendation at iteration $t$, $T_{\text{train}}^{(t)}$ is the training time for the recommended configuration, and $T_{\text{eval}}^{(t)}$ is the validation evaluation time.
Computing each technique separately enables direct comparison of computational efficiency between Bayesian Optimization and LLM-guided optimization, facilitating assessment of the trade-off between recommendation quality and computational cost.
4.2 Results and Discussion
@@ 4.2.1 RQ1 (Performance)
Figure 2: Validation RMSE convergence across optimization strategies. LLM-AutoOpt (blue) achieves the lowest final RMSE and fastest convergence compared to BO-only (orange) and LLM without meta-knowledge (green) approaches.
Table 5: Comparison of Mean RMSE, Average Training Time, and Optimization Time for different methods (with standard deviations).
@@ 4.2.4 RQ4: Impact of Meta-Knowledge on Optimization Quality
Table 6: Hyperparameter configurations and RMSE results across optimization trials with method comparisons.
@@
4.2.5 RQ5: Interpretability and Reasoning Transparency
@@
Overall, incorporating meta-knowledge into LLM-AutoOpt leads to (i) improved forecasting accuracy and faster convergence compared with manual, Bayesian, and meta-agnostic LLM optimization methods, (ii) enhanced interpretability through structured priors and transparent reasoning that connect dataset characteristics to hyperparameter decisions, and (iii) a trade-off between performance and computational cost, with effectiveness depending on prompt design and requiring broader evaluation to confirm scalability and generalization.